OpenStudy (rainbow_rocks03):

PLZ HELP...medal+fan, NO DIRECT ANSWERS. Meredith wants to paint 3 of the walls in her room.Each wall with dimensions 10ft by 12ft.She has 1 can of paint that will cover approximately 420 square ft.Which method can Meredith use to determine whether or not she has enough paint to complete the job? A.multiply 10 by 12.compare the product to 420 B.multiply 10 by 12 and divide by 3.compare the result to 420. C.multiply 10 by 12 by 3.compare the product to 420. D.multiply 10 by 12 by 4 and divide by 3.compare the result to 420.

OpenStudy (rainbow_rocks03):

@OM14forever @texaschic101

OpenStudy (rainbow_rocks03):

@Nnesha @sammixboo

OpenStudy (rainbow_rocks03):

@Hero @whpalmer4

OpenStudy (anonymous):

10*12=?

OpenStudy (rainbow_rocks03):

10*12=120

OpenStudy (anonymous):

10*12=120 Since there are 3 walls in the room..... 120*3=??????

OpenStudy (rainbow_rocks03):

120*3=360

OpenStudy (anonymous):

And she has enough paint to cover 420 square ft

OpenStudy (anonymous):

So C is the answer

OpenStudy (rainbow_rocks03):

TY!